Solve the differential equation given by

(3x + 2y + y^2)dx + (x+4xy+5y^2)dy=0


1
Expert's answer
2021-08-26T17:01:55-0400
"\dfrac{\partial Q}{\partial x}=1+4y, \dfrac{\partial P}{\partial y}=2+2y""\dfrac{\partial Q}{\partial x}\not=\dfrac{\partial P}{\partial y}""\mu=x+y^2""(x+y^2)(3x + 2y + y^2)dx""+ (x+y^2)(x+4xy+5y^2)dy=0""P=3x^2+2xy+xy^2+3xy^2+2y^3+y^4""\dfrac{\partial P}{\partial y}=2x+8xy+6y^2+4y^3""Q=x^2+4x^2y+5xy^2+xy^2+4xy^3+5y^4""\dfrac{\partial Q}{\partial x}=2x+8xy+6y^2+4y^3""\dfrac{\partial P}{\partial y}=\dfrac{\partial Q}{\partial x}""\dfrac{\partial u}{\partial x}=3x^2+2xy+4xy^2+2y^3+y^4"

Integrate



"u=x^3+x^2y+2x^2y^2+2xy^3+xy^4+\varphi(y)"

Then


"\dfrac{\partial u}{\partial y}=x^2+4x^2y+6xy^2+4xy^3+\varphi'(y)""=Q=x^2+4x^2y+6xy^2+4xy^3+5y^4""\varphi'(y)=5y^4""\varphi(y)=y^5+C"

The original differential equation has the following solution



"x^3+x^2y+2x^2y^2+2xy^3+xy^4-y^5=C"